What is the algebra fun solution to prove that 1 + 1 = 3?
Found 2 solutions by mukhopadhyay, AnlytcPhil:
Answer by mukhopadhyay(490)

The funny part is: 1 + 1 is never 3; where did you get this question from?

Answer by AnlytcPhil(1806)

What is the algebra fun solution to prove that 1 + 1 = 3?

Start with:

(1)          x² + 2x² = 3x²

Let y = x, then x² = y² = xy
So, in (1), leave the first x² as it is,
replace the second x² by y²
and replace the third x² by xy

(2)          x² + 2y² = 3xy

Subtract 3y² from both sides

       x² + 2y² - 3y² = 3xy - 3y²

              x² - y² = 3xy - 3y²

Factor both sides

       (x - y)(x + y) = 3y(x - y)

Cancel the (x - y)'s

                x + y = 3y

Since x = y, substitute y for the x

                y + y = 3y

Divide every term by y

                1 + 1 = 3

Can you find the flaw?

Edwin
AnlytcPhil@aol.com